Fermi National Laboratory
 

Accelerator Update for November 28, 2007 - November 30, 2007

 

Wednesday November 28

The day shift began with the Tevatron (TeV) F sector in access, with Operations monitoring the Recycler stash (242.2E10), with the Antiproton Source (Pbar) stacking (171.6E10), with the Main Injector (MI), Booster, and Linac operating normally, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) waiting for beam, and with the NuMI and SciBooNE / MiniBooNE experiments off due to the TeV F sector access.

Operations prepared the TeV sectors A-E for access at 8:49 AM.

FESS technicians reported at 8:53 AM that they were beginning their repair of the B0 pond water pump.

CDF went in on access at 10:08 AM.

Preaccelerator experts accessed the I- Source dome at 10:21 AM. They reported that the generator brushes needed to be changed and the hydrogen bottle needed to be replaced.

Controls experts reported at 11:04 AM on the clock and data logger problems from Tuesday in the P1/P2 transfer lines. They found that the C290 cards had lost their clock sync with the Front-End. A reset fixed the problem and it appeared that only the MI-1 Front End was affected.

Vacuum system techs came out of F sector. They reported that they had found and fixed a vacuum leak just upstream of the F11 beam valve. Operations resumed sending beam to Sci/MiniBooNE at 12:01 PM.

Operations resumed sending beam to Pbar and NuMI at 12:46 PM. Pbar experts began a Pbar Debuncher study.

CDF came out of access at 1:20 PM.

A MI expert began a slipstacking study at 4:31 PM.

Operations put the TeV through a QBS ramp at 4:03 PM.

FESS techs reported at 4:11 PM that they had completed their work on the B0 pond pump.

At 4:29 PM, all the Linac Klystron RF stations tripped off due to a common LCW water skid trip. The cause of the trip was uncertain. Linac experts investigated.

Pbar experts completed their study at 5:13 PM.

Operations put the TeV through its first high-energy ramp at 6:05 PM.

At 6:19 PM, a D0 quadrupole power supply (D0Q5) tripped off on a lead failure indication. A TeV expert reported that the lead tolerances were set wrong.

At 6:22 PM, KRF7 tripped off. It reset.

Operations put the TeV through a dry squeeze at 6:24 PM.

At 7:04 PM, the TeV suffered a quench at sector F2. It appeared that the kickers did not fire due to a clock problem. Operators contacted a controls expert and then went with a TeV expert to reset the sector F17 465 card locally.

Operations put the TeV through a QBS ramp at 8:24 PM, through a high-energy ramp at 8:32 PM, and through a dry squeeze study at 8:43 PM.

TeV experts finished their study. Operations prepared D0 for an access at 11:22 PM.

Thursday November 29

The midnight shift began with D0 in access, with Operations monitoring the stash (238.0E10), with Pbar stacking (175.2E10), and with NuMI and Sci /MiniBooNE taking beam.

At 12:01 AM, Operations reported that KRF7 tripped off repeatedly on a reflected power problem. Operators investigated.

At 12:21 AM, Operations reported that the TeV BPMs were not reporting. Operations contacted experts.

At 12:35 AM, KRF7 tripped off for the first of many times in about ten minutes. It reset.

At 1:17 AM, KRF7 tripped off and would not reset. Operations called in experts.

D0 came out of access at 1:34 AM.

Operations put the TeV through a dry squeeze at 2:23 AM.

At 3 AM, Operations reported that the Pbar stack had started unstacking when Linac went off. Operations contacted an expert and then adjusted the cooling for its core. About 15E10 antiprotons were lost.

Operations reported at 3:55 PM that they Recycler had stashed antiprotons twice during the day shift.

Linac experts reported at 4:41 AM that they had fixed KRF7. Operations resumed sending beam to Pbar, NuMI, and Sci/MiniBooNE.

Operations put the TeV through a wet squeeze at 4:55 AM.

At 5:01 AM, the TeV suffered an abort due to a sector F1 BLM. Operators reported that some F17 collimators had gone into alarm when the abort was pulled. Their investigations continued.

At 5:35 AM, Operations ramped the TEV to 150 GeV, unmasked the BLMs, and a minute later the TeV aborted. This time the abort appeared to be caused by a corrector magnet (V714) power supply timing glitch. There was also an indication of some poor quality beam coming out of MI.

Operation began tuning the TeV at 7:12 AM for a proton only study. During this tuning operators talked with experts about the F17 collimators. Apparently, the collimators were acting oddly due to another clock problem. A controls expert was called in to investigate.

At 9:15 AM, a controls expert reported that he found and fixed a bad time table for the F17 kickers.

Operations put the TeV into shot setup at 10:30 AM.

Operations turned off the NuMI beam and prepared the area for access at 11 AM.

Pbar experts began a 15-minute target scan at 12:47 PM.

Operations established store 5760 at 1:06 PM with an initial luminosity of 166.82E30.

Preaccelerator experts reported at 3:30 PM that they had replaced the generator brushes and installed a new hydrogen bottle for the I- Source.

Operation resumed sending beam to Pbar and Sci/MiniBooNE at 3:35 PM.

At 9:26 PM, Pbar set another stacking record with a rate of 23.87 mA per hour.

Operations reported at midnight that they Recycler had stashed twice during the evening shift.

Friday November 30

The midnight shift began with Operations monitoring store 5760 (50.71E30) and a stash (261.1E10), with Pbar stacking (97.7E10), with NuMI in access, and with Sci/MiniBooNE taking beam.

NuMI came out of access at 1:29 AM and Operations established beam to the experiment at 1:55 AM.

Operations terminated store 5760 at 2:06 AM.

Operations prepared D0 for an access at 2:11 AM. They came out of access at 2:28 AM.

Operations put the TeV into shot setup at 2:36 AM.

Operations established store 5761 at 4:45 AM with an initial luminosity of 172.6E30.

Operations reported at 7:59 AM that they Recycler had stashed twice during the midnight shift.

The Plans for Friday and the Weekend

The plans for today and the weekend are to stack, stash, and store. Later today, the Run Coordinator will allow D0 and CDF a two to four hour access to make necessary repairs on their detectors.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 11/30/2007 -->    email Fermilab